OpenStudy (anonymous):

Let f(x)=x^2-3x and P=(1, f(1)). Find the slope of the curve y=f(x) at P

OpenStudy (anonymous):

slope=2x-3=2*1-3=-1

OpenStudy (anonymous):

how is the slope 2x-3?

OpenStudy (anonymous):

first derivative of a function at a point gives the slope at that point

OpenStudy (anonymous):

what is the first derivative in this case?